OpenStudy (anonymous):

The equation of a line is 15y + 8x = 19. a. Find the x-intercept. b. Find the y-intercept.

OpenStudy (campbell_st):

to find the x intercept substitute y = 0 and solve for x to find the y intercept, substitute x = 0 into the original equation and solve for y

OpenStudy (anonymous):

I don't quite understand how to do that wouldn't it be 15y+8(0) = 19 and then 15 + 0 = 19, what did I do wrong?

OpenStudy (campbell_st):

no thats correct for the y intercept... so you need to solve 15y = 19 so y = 19/15 so the y intercept is (0, 19/15) similar process for the x- intercept